# Explainability-Based Token Replacement on LLM-Generated Text

### v1

@@ 4. Methodology
We built an ensemble of multiple transformers instead of relying on a single one to classify human and AIGT), as inspired by Mohammadi et al. (2023). In this method, we use the strengths of each model and create an ensemble that can be fine-tuned on similar datasets. This helps the model adapt to different domains and reduces problems caused by limited data, as the features learned by each part complement each other. Specifically, we integrated bert-base-multilingual-uncased, distilbert-base-multilingual-cased, and xlm-roberta-base. Each transformer produced a representation of the input text, and we concatenated these representations before passing them through a dense layer with L2 regularization for final classification.
We first built a Combined BERT model by training three BERT models on the AuTexTification training set. We optimized the network’s hyperparameters and evaluated performance on the AuTexTification test set, focusing on F1 and average metrics across folds. After training, we froze these model weights to retain the features they had learned and reduce the risk of overfitting. We next created a combined model that merges three BERT models with frozen weights (trained on AuTexTification) and three fresh BERT models with unfrozen parameters. These fresh models were fine-tuned on our newly augmented dataset Such an ensemble synergy aligns with Fraser et al. (2025), who highlight the benefits of multi-model integration in detecting AI-generated text. Algorithm 1 outlines the training and process for our combined BERT ensemble. During training, we minimize the binary cross-entropy loss on each branch using the augmented data. We then unify the outputs in a final fully connected layer, optionally employing majority voting if multiple final heads are used. Figure 1 shows the architecture, with frozen models on the left (gray) and fresh models on the right (green).
Figure 1. The architecture of the proposed ensemble model. Outputs of three frozen BERT models (left) and three fresh BERT models (right) are concatenated and passed through dense layers, followed by a voting mechanism.
@@ 4.1. Experimental Setup
The final training stage used 90% of the augmented development set, while we left 10% for test. The hyperparameters and early-stopping rules were adjusted using the validation split. After training, we evaluated the system on the remaining data, measuring precision, recall, F1-score, and accuracy. By combining different transformers, freezing selectively learned features, and training fresh models in parallel, this ensemble method balanced the advantages of robust pre-trained representations with the adaptability needed for the new dataset.
@@ 4.2. Explainability-driven detection
We follow the methodology in (Mohammadi et al., 2024), in this framework, we only analyze sentences that the model correctly classifies, ensuring the tokens we highlight have genuinely contributed to accurate decisions. We then average each token’s importance within the sentence to get a consistent measure of its overall effect. This step helps us confirm that the extracted terms are meaningful indicators of the model’s reasoning. Given a trained model $f$ and an instance $x$ composed of tokens $(t_{1},t_{2},\ldots,t_{m})$, SHAP assigns a Shapley value $\phi_{j}$ to each token $t_{j}$. These values are computed by measuring how much including or excluding $t_{j}$ changes the model’s output, across all possible subsets of tokens. Formally:
Here, $f(S)$ is the model’s prediction when only the tokens in set $S$ are present. The weighting term ensures each token’s effect is fairly distributed among all subsets (Lundberg and Lee, 2017). LIME, on the other hand, makes small changes near the original text to learn a simpler, local model that approximates $f$. It then assigns coefficients reflecting each token’s importance for that instance (Ribeiro et al., 2016). Although LIME focuses more on local neighborhoods, it also provides per-token contributions based on how the model’s prediction shifts when individual tokens are perturbed. We also consider a random token selection. This lets us compare how much improvement we gain by targeting the tokens that SHAP or LIME indicates are most influential. By using both a global (SHAP) and local (LIME) perspective, we can better trust that the identified tokens genuinely drive the model’s decision. In our subsequent steps, we use these tokens to guide targeted replacements and measure how effectively such edits can mask AI-generated content.
@@ 4.4. Evaluation Metrics
We use a combination of detection and text-similarity metrics to evaluate how rewriting strategies affect our models’ ability to detect AIGT, as well as how much the rewritten text deviates from its original form. First, once we rewrite an AIGT via HSR, PSR, GPT, or GPT+Genre. Next, we investigate how different rewriting scenarios affect AI-text detection. We adopt the adversarial rewriting perspective discussed in Fraser et al. (2025), where paraphrased or partially edited texts can undermine naive detectors. We run it through each classifier to check whether the label remains AI or flips to human. We then calculate precision, recall, F1, and accuracy to capture each model’s classification performance to detect AI/human texts under these modified inputs. Additionally, we compute the flip rate, defined as the fraction of AI texts whose predicted label changes to human after rewriting; a higher flip rate implies that the rewriting method is more effective in evading detection. In parallel, we calculate the degree of textual change introduced by each rewriting approach. We use BLEU, ROUGE-1, and ROUGE-L to measure the overlap between the original text and its rewritten form, with higher scores showing fewer modifications. By combining detection metrics, and text-similarity scores, we gain a clear view of how effectively each rewriting method hides AI-generated content and how much it changes the original text.
@@ 5. Experimental Analysis and Results
In this section, we compare five classification models (XGBoost, BERT-base, DistilBERT, XLM-RoBERTa, and our Ensemble approach) across various token-replacement scenarios for the detection of AIGT, Consistent with Fraser et al. (2025), we anticipate that multi-domain evaluation is critical for robust results. Our main objective is to observe how different rewriting methods (HSR, PSR, GPT, GPT+Genre) combined with token-selection strategies (SHAP, LIME, Random) affect each model’s ability to detect generated text.
Table 3 shows each model’s performance on unmodified AIGT across two languages (English, Dutch) and three domains (news, reviews, and X/Twitter). As can be seen, although certain single models perform well on specific subsets (e.g., DistilBERT in English news, XGBoost in English Twitter), the Ensemble approach tends to achieve the best performance in most cases. This pattern is particularly clear in Dutch news, where the Ensemble model attains perfect scores. Even when other models have high recall or precision on certain tasks, they often fall behind on at least one metric (e.g., XLM-RoBERTa predicts too many positives, increasing recall but lowering precision). showing strong cross-language adaptability. This result agrees with  Fraser et al. (2025), who emphasize the importance of multi-lingual calibration for robust detection.
Table 3. Model performance by language and domain on unmodified AI text. Bold indicates the best result within each language and domain block.
@@
Table 4. F1 score and Accuracy of each detector under every rewriting scenario. The best metric in a column is bolded; the largest negative drop is -00%, the smallest negative drop is -00%. Δ=(Scenario ValueOriginal $\Delta=(\text{Scenario Value}-\text{Original Value})\times 100\%$.
@@
F1 ($\Delta$) Acc ($\Delta$) F1 ($\Delta$) Acc ($\Delta$) F1 ($\Delta$) Acc ($\Delta$) F1 ($\Delta$) Acc ($\Delta$) F1 ($\Delta$) Acc ($\Delta$)
@@
SHAP HSR 0.73 (-8%) 0.75 (-5%) 0.71 (-3%) 0.72 (-3%) 0.76 (-3%) 0.75 (-2%) 0.68 (0%) 0.54 (0%) 0.78 (-5%) 0.80 (-3%)
SHAP PSR 0.78 (-3%) 0.78 (-2%) 0.73 (-1%) 0.74 (-1%) 0.78 (-1%) 0.77 (0%) 0.68 (0%) 0.54 (0%) 0.82 (-1%) 0.82 (-1%)
SHAP GPT 0.79 (-2%) 0.79 (-1%) 0.73 (-1%) 0.74 (-1%) 0.77 (-2%) 0.76 (-1%) 0.68 (0%) 0.55 (+1%) 0.80 (-3%) 0.81 (-2%)
SHAP GPT+Genre 0.80 (-1%) 0.80 (0%) 0.73 (-1%) 0.74 (-1%) 0.77 (-2%) 0.76 (-1%) 0.68 (0%) 0.55 (+1%) 0.79 (-4%) 0.80 (-3%)
LIME HSR 0.25 (-56%) 0.48 (-32%) 0.56 (-18%) 0.62 (-13%) 0.62 (-17%) 0.65 (-12%) 0.67 (-1%) 0.52 (-2%) 0.64 (-19%) 0.69 (-14%)
LIME PSR 0.56 (-25%) 0.63 (-17%) 0.64 (-10%) 0.67 (-8%) 0.71 (-8%) 0.71 (-6%) 0.66 (-2%) 0.52 (-2%) 0.73 (-10%) 0.75 (-8%)
LIME GPT 0.57 (-24%) 0.64 (-16%) 0.73 (-1%) 0.74 (-1%) 0.78 (-1%) 0.77 (0%) 0.68 (0%) 0.55 (+1%) 0.83 (0%) 0.83 (0%)
LIME GPT+Genre 0.60 (-21%) 0.65 (-15%) 0.75 (+1%) 0.75 (0%) 0.77 (-2%) 0.76 (-1%) 0.68 (0%) 0.55 (+1%) 0.81 (-2%) 0.81 (-2%)
Random HSR 0.73 (-8%) 0.74 (-6%) 0.63 (-11%) 0.67 (-8%) 0.68 (-11%) 0.69 (-8%) 0.67 (-1%) 0.53 (-1%) 0.73 (-10%) 0.75 (-8%)
Random PSR 0.76 (-5%) 0.76 (-4%) 0.65 (-9%) 0.68 (-7%) 0.74 (-5%) 0.73 (-4%) 0.67 (-1%) 0.53 (-1%) 0.78 (-5%) 0.79 (-4%)
Random GPT 0.74 (-7%) 0.75 (-5%) 0.74 (0%) 0.75 (0%) 0.77 (-2%) 0.76 (-1%) 0.68 (0%) 0.55 (+1%) 0.81 (-2%) 0.82 (-1%)
Random GPT+Genre 0.75 (-6%) 0.75 (-5%) 0.73 (-1%) 0.74 (-1%) 0.78 (-1%) 0.77 (0%) 0.68 (0%) 0.55 (+1%) 0.81 (-2%) 0.81 (-2%)
@@
HSR and PSR replace tokens with synonyms from a human corpus, often preserving the original text’s structure. This is reflected in relatively high BLEU scores (e.g., over 90% in the SHAP+HSR or SHAP+PSR rows), meaning minimal rewriting. Nonetheless, these small edits can still mislead detectors if the replaced tokens were crucial. GPT-based rewrites typically lower BLEU into the 70s or 80s, reflecting more substantial paraphrases. In turn, many detectors (e.g., XGBoost or XLM-RoBERTa) experience an increase in recall but a decrease in precision, or the opposite, because the text shifts away from familiar AI phrasing. GPT+Genre further contextualizes the rewrite with domain hints (e.g. write it like a news article), causing domain-sensitive lexical changes. This can introduce new phrasing that confuses detectors reliant on domain-specific cues.
@@
XLM-RoBERTa often shows high recall (sometimes hitting 1.0), but its precision remains around 0.52, leading to only modest changes in F1. DistilBERT is more balanced yet remains vulnerable when its top keywords are replaced by synonyms or paraphrases. BERT-base typically holds a moderate advantage across scenarios but can be undermined in LIME-based rewrites (e.g., LIME+HSR, where BERT-base F1 falls to about 0.56). By contrast, the Ensemble approach consistently retains the highest or near-highest F1 and accuracy. For example, under LIME+GPT, the Ensemble still attains an F1 of 0.83, significantly above XGBoost’s 0.57 or BERT-base’s 0.73. Similarly, in SHAP plus GPT+Genre, the Ensemble preserves a strong F1 of about 0.79, while other models vary more dramatically. Also, We show text similarity scores (BLEU, ROUGE1, ROUGEL) to capture how extensively each rewrite modifies the original text.
@@
Table 5 presents text similarity scores (BLEU, ROUGE1, ROUGEL) under all these rewriting conditions. Looking at BLEU and ROUGE scores confirms that GPT-based methods, especially GPT+Genre, introduce the largest textual shifts. Replacements guided by synonyms (HSR or PSR) keep BLEU near or above 90, whereas GPT rewrites can dip into the mid-70s. This trade-off highlights how more extensive rewriting can more thoroughly remove AI signatures at the expense of text fidelity.
@@
NoRewriting Original 1,00 1,00 1,00
SHAP HSR 0,93 (-7%) 0,97 (-3%) 0,97 (-3%)
SHAP PSR 0,97 (-3%) 0,99 (-1%) 0,99 (-1%)
SHAP GPT 0,86 (-14%) 0,93 (-7%) 0,93 (-7%)
SHAP GPT+Genre 0,86 (-14%) 0,94 (-6%) 0,94 (-6%)
LIME HSR 0,84 (-16%) 0,92 (-8%) 0,92 (-8%)
LIME PSR 0,90 (-10%) 0,96 (-5%) 0,95 (-5%)
LIME GPT 0,76 (-24%) 0,84 (-16%) 0,84 (-16%)
LIME GPT+Genre 0,76 (-24%) 0,85 (-15%) 0,84 (-16%)
Random HSR 0,84 (-16%) 0,92 (-8%) 0,92 (-8%)
Random PSR 0,90 (-11%) 0,95 (-5%) 0,95 (-5%)
Random GPT 0,76 (-24%) 0,84 (-16%) 0,84 (-16%)
Random GPT+Genre 0,76 (-24%) 0,84 (-16%) 0,84 (-16%)
@@
BERT $120$ $7$ $5.83$
DistilBERT $120$ $9$ $7.50$
XLM-R $120$ $1$ $0.83$
Ensemble $120$ $8$ $6.67$
PSR XGBoost $120$ $7$ $5.83$
BERT $120$ $4$ $3.33$
DistilBERT $120$ $4$ $3.33$
XLM-R $120$ 00 $0.00$
Ensemble $120$ $3$ $2.50$
@@
BERT $120$ $20$ $16.67$
DistilBERT $120$ $19$ $15.83$
XLM-R $120$ $5$ $4.17$
Ensemble $120$ $19$ $15.83$
@@
SHAP GPT XGBoost $120$ $4$ $3.33$
BERT $120$ $7$ $5.83$
DistilBERT $120$ $7$ $5.83$
XLM-R $120$ 00 $0.00$
@@
GPT+Genre XGBoost $120$ $4$ $3.33$
BERT $120$ $3$ $2.50$
DistilBERT $120$ $6$ $5.00$
XLM-R $120$ 00 $0.00$
@@
BERT $120$ $4$ $3.33$
DistilBERT $120$ $7$ $5.83$
XLM-R $120$ 00 $0.00$
Ensemble $120$ $9$ $7.50$
@@
BERT $120$ $6$ $5.00$
DistilBERT $120$ $6$ $5.00$
XLM-R $120$ 00 $0.00$
Ensemble $120$ $5$ $4.17$
@@
As shown in Table 6, many rewriting strategies cause a lot of flips for some detectors, especially the simpler ones. However, the Ensemble model usually has a lower flip rate. This shows that using several models together makes detection more reliable. Figure 2 also shows how the flipped items overlap for each rewriting method. Even when many detectors are affected by these edits, the Ensemble is less likely to be fooled, highlighting the advantage of using an ensemble approach

### v2

@@ 4. Methodology
We built an ensemble of multiple transformers instead of relying on a single one to classify human and AIGT), as inspired by mohammadi2023towards. In this method, we use the strengths of each model and create an ensemble that can be fine-tuned on similar datasets. This helps the model adapt to different domains and reduces problems caused by limited data, as the features learned by each part complement each other. Specifically, we integrated bert-base-multilingual-uncased, distilbert-base-multilingual-cased, and xlm-roberta-base. Each transformer produced a representation of the input text, and we concatenated these representations before passing them through a dense layer with L2 regularization for final classification.
We first built a Combined BERT model by training three BERT models on the AuTexTification training set. We optimized the network’s hyperparameters and evaluated performance on the AuTexTification test set, focusing on F1 and average metrics across folds. After training, we froze these model weights to retain the features they had learned and reduce the risk of overfitting. We next created a combined model that merges three BERT models with frozen weights (trained on AuTexTification) and three fresh BERT models with unfrozen parameters. These fresh models were fine-tuned on our newly augmented dataset Such an ensemble synergy aligns with fraser2025detecting, who highlight the benefits of multi-model integration in detecting AI-generated text. Algorithm 1 outlines the training and process for our combined BERT ensemble. During training, we minimize the binary cross-entropy loss on each branch using the augmented data. We then unify the outputs in a final fully connected layer with L2 regularization. Figure 1 shows the architecture, with frozen models on the left (gray) and fresh models on the right (green).
Figure 1. The architecture of the proposed ensemble model. Outputs of three frozen BERT models (left) and three fresh BERT models (right) are concatenated and passed through a dense layer followed by a sigmoid activation (and threshold at 0.5) for final binary predictions.
@@ 4.1. Experimental Setup
We trained on the augmented training split and selected hyperparameters using a validation split drawn from training data. Final evaluation was performed on a held-out non-augmented CLIN33 test set of 240 texts (120 human, 120 AI), stratified across two languages and three domains (40 texts per domain per language). This same fixed test set was used for all experiments to ensure fair comparisons. The hyperparameters and early-stopping rules were adjusted using the validation split. After training, we evaluated the system on the remaining data, measuring precision, recall, F1-score, and accuracy. By combining different transformers, freezing selectively learned features, and training fresh models in parallel, this ensemble method balanced the advantages of robust pre-trained representations with the adaptability needed for the new dataset.
Computational Resources.
Table 2 summarizes the computational requirements of each model in our experiments. All training was conducted on NVIDIA A100 GPUs (40GB memory). The ensemble model, while requiring substantially more resources than individual models, remains tractable for research settings and can be parallelized across multiple GPUs to reduce wall-clock training time. For inference, the ensemble processes samples sequentially through each component, with the total inference time being the sum of individual model times plus the final dense layer computation. In production scenarios, inference latency could be reduced through model distillation or parallel execution of the ensemble components.
Table 2. Computational cost comparison of detection models. Training times are approximate and based on the CLIN33 dataset. Inference times represent average per-sample latency on GPU. Frozen branch $\approx$454M parameters; fresh branch $\approx$454M parameters; total parameters loaded for the ensemble $\approx$908M (+ classifier head).
Model Parameters Training Time Inference (ms) GPU Memory
XGBoost $\sim$50K $\sim$5 min $\sim$1 CPU only
BERT-base 110M $\sim$2 h $\sim$50 $\sim$4 GB
DistilBERT 66M $\sim$1.5 h $\sim$30 $\sim$2 GB
XLM-RoBERTa 278M $\sim$3 h $\sim$60 $\sim$6 GB
Ensemble (ours) $\sim$908M $\sim$8 h total $\sim$200 $\sim$16 GB
@@ 4.2. Explainability-driven detection
We follow the methodology in (mohammadi2024transparent), in this framework, we only analyze sentences that the model correctly classifies, ensuring the tokens we highlight have genuinely contributed to accurate decisions. We then average each token’s importance within the sentence to get a consistent measure of its overall effect. This step helps us confirm that the extracted terms are meaningful indicators of the model’s reasoning. Given a trained model $f$ and an instance $x$ composed of tokens $(t_{1},t_{2},\ldots,t_{m})$, SHAP assigns a Shapley value $\phi_{j}$ to each token $t_{j}$. These values are computed by measuring how much including or excluding $t_{j}$ changes the model’s output, across all possible subsets of tokens. Formally:
Here, $f(S)$ is the model’s prediction when only the tokens in set $S$ are present. The weighting term ensures each token’s effect is fairly distributed among all subsets (lundberg2017unified). LIME, on the other hand, makes small changes near the original text to learn a simpler, local model that approximates $f$. It then assigns coefficients reflecting each token’s importance for that instance (ribeiro2016should). Although LIME focuses more on local neighborhoods, it also provides per-token contributions based on how the model’s prediction shifts when individual tokens are perturbed. We also consider a random token selection. This lets us compare how much improvement we gain by targeting the tokens that SHAP or LIME indicates are most influential. By using both a global (SHAP) and local (LIME) perspective, we can better trust that the identified tokens genuinely drive the model’s decision. In our subsequent steps, we use these tokens to guide targeted replacements and measure how effectively such edits can mask AI-generated content.
@@ 4.4. Evaluation Metrics
We use a combination of detection and text-similarity metrics to evaluate how rewriting strategies affect our models’ ability to detect AIGT, as well as how much the rewritten text deviates from its original form. First, once we rewrite an AIGT via HSR, PSR, GPT, or GPT+Genre. Next, we investigate how different rewriting scenarios affect AI-text detection. We adopt the adversarial rewriting perspective discussed in fraser2025detecting, where paraphrased or partially edited texts can undermine naive detectors. We run it through each classifier to check whether the label remains AI or flips to human. We then calculate precision, recall, F1, and accuracy to capture each model’s classification performance to detect AI/human texts under these modified inputs. Additionally, we compute the flip rate, defined as the fraction of AI texts whose predicted label changes to human after rewriting; a higher flip rate implies that the rewriting method is more effective in evading detection. In parallel, we calculate the degree of textual change introduced by each rewriting approach. We use BLEU, ROUGE-1, and ROUGE-L to measure the overlap between the original text and its rewritten form, with higher scores showing fewer modifications. By combining detection metrics, and text-similarity scores, we gain a clear view of how effectively each rewriting method hides AI-generated content and how much it changes the original text.
@@ 5. Experimental Analysis and Results
In this section, we compare five classification models (XGBoost, BERT-base, DistilBERT, XLM-RoBERTa, and our Ensemble approach) across various token-replacement scenarios for the detection of AIGT, Consistent with fraser2025detecting, we anticipate that multi-domain evaluation is critical for robust results. Our main objective is to observe how different rewriting methods (HSR, PSR, GPT, GPT+Genre) combined with token-selection strategies (SHAP, LIME, Random) affect each model’s ability to detect generated text.
Table 4 shows each model’s performance on unmodified AIGT across two languages (English, Dutch) and three domains (news, reviews, and X/Twitter). As can be seen, although certain single models perform well on specific subsets (e.g., DistilBERT in English news, XGBoost in English Twitter), the Ensemble approach tends to achieve the best performance in most cases. This pattern is particularly clear in Dutch news, where the Ensemble model achieves near-perfect scores (F1=0.99, Precision=0.98, Recall=1.00). Even when other models have high recall or precision on certain tasks, they often fall behind on at least one metric (e.g., XLM-RoBERTa predicts too many positives, increasing recall but lowering precision). showing strong cross-language adaptability. This result agrees with  fraser2025detecting, who emphasize the importance of multi-lingual calibration for robust detection.
Table 4. Model performance by language and domain on the original (unmodified) test set. Bold indicates the best result within each language and domain block.
@@
English Average Dutch Average
XGBoost avg 0.77 0.81 0.80 0.80 XGBoost avg 0.73 0.75 0.71 0.72
BERT-base avg 0.78 0.79 0.84 0.81 BERT-base avg 0.63 0.66 0.55 0.60
DistilBERT avg 0.82 0.81 0.88 0.84 DistilBERT avg 0.65 0.66 0.65 0.65
XLM-RoBERTa avg 0.55 0.53 0.97 0.68 XLM-RoBERTa avg 0.53 0.52 0.83 0.63
Ensemble avg 0.85 0.86 0.86 0.86 Ensemble avg 0.82 0.80 0.76 0.78
@@
[…]
@@
F1 ($\Delta$ $p$) Acc ($\Delta$ $p$) F1 ($\Delta$ $p$) Acc ($\Delta$ $p$) F1 ($\Delta$ $p$) Acc ($\Delta$ $p$) F1 ($\Delta$ $p$) Acc ($\Delta$ $p$) F1 ($\Delta$ $p$) Acc ($\Delta$ $p$)
@@
SHAP HSR 0.73 (-8 pp) 0.75 (-5 pp) 0.71 (-3 pp) 0.72 (-3 pp) 0.76 (-3 pp) 0.75 (-2 pp) 0.68 (0 pp) 0.54 (0 pp) 0.78 (-5 pp) 0.80 (-3 pp)
SHAP PSR 0.78 (-3 pp) 0.78 (-2 pp) 0.73 (-1 pp) 0.74 (-1 pp) 0.78 (-1 pp) 0.77 (0 pp) 0.68 (0 pp) 0.54 (0 pp) 0.82 (-1 pp) 0.82 (-1 pp)
SHAP GPT 0.79 (-2 pp) 0.79 (-1 pp) 0.73 (-1 pp) 0.74 (-1 pp) 0.77 (-2 pp) 0.76 (-1 pp) 0.68 (0 pp) 0.55 (+1 pp) 0.80 (-3 pp) 0.81 (-2 pp)
SHAP GPT+Genre 0.80 (-1 pp) 0.80 (0 pp) 0.73 (-1 pp) 0.74 (-1 pp) 0.77 (-2 pp) 0.76 (-1 pp) 0.68 (0 pp) 0.55 (+1 pp) 0.79 (-4 pp) 0.80 (-3 pp)
LIME HSR 0.25 (-56 pp) 0.48 (-32 pp) 0.56 (-18 pp) 0.62 (-13 pp) 0.62 (-17 pp) 0.65 (-12 pp) 0.67 (-1 pp) 0.52 (-2 pp) 0.64 (-19 pp) 0.69 (-14 pp)
LIME PSR 0.56 (-25 pp) 0.63 (-17 pp) 0.64 (-10 pp) 0.67 (-8 pp) 0.71 (-8 pp) 0.71 (-6 pp) 0.66 (-2 pp) 0.52 (-2 pp) 0.73 (-10 pp) 0.75 (-8 pp)
LIME GPT 0.57 (-24 pp) 0.64 (-16 pp) 0.73 (-1 pp) 0.74 (-1 pp) 0.78 (-1 pp) 0.77 (0 pp) 0.68 (0 pp) 0.55 (+1 pp) 0.83 (0 pp) 0.83 (0 pp)
LIME GPT+Genre 0.60 (-21 pp) 0.65 (-15 pp) 0.75 (+1 pp) 0.75 (0 pp) 0.77 (-2 pp) 0.76 (-1 pp) 0.68 (0 pp) 0.55 (+1 pp) 0.81 (-2 pp) 0.81 (-2 pp)
Random HSR 0.73 (-8 pp) 0.74 (-6 pp) 0.63 (-11 pp) 0.67 (-8 pp) 0.68 (-11 pp) 0.69 (-8 pp) 0.67 (-1 pp) 0.53 (-1 pp) 0.73 (-10 pp) 0.75 (-8 pp)
Random PSR 0.76 (-5 pp) 0.76 (-4 pp) 0.65 (-9 pp) 0.68 (-7 pp) 0.74 (-5 pp) 0.73 (-4 pp) 0.67 (-1 pp) 0.53 (-1 pp) 0.78 (-5 pp) 0.79 (-4 pp)
Random GPT 0.74 (-7 pp) 0.75 (-5 pp) 0.74 (0 pp) 0.75 (0 pp) 0.77 (-2 pp) 0.76 (-1 pp) 0.68 (0 pp) 0.55 (+1 pp) 0.81 (-2 pp) 0.82 (-1 pp)
Random GPT+Genre 0.75 (-6 pp) 0.75 (-5 pp) 0.73 (-1 pp) 0.74 (-1 pp) 0.78 (-1 pp) 0.77 (0 pp) 0.68 (0 pp) 0.55 (+1 pp) 0.81 (-2 pp) 0.81 (-2 pp)
@@
HSR and PSR replace tokens with synonyms from a human corpus, often preserving the original text’s structure. This is reflected in relatively high BLEU scores (e.g., over 90% in the SHAP+HSR or SHAP+PSR rows), meaning minimal rewriting. Nonetheless, these small edits can still mislead detectors if the replaced tokens were crucial. GPT-based rewrites typically lower BLEU scores to the range of 0.76–0.86, reflecting more substantial paraphrases. In turn, many detectors (e.g., XGBoost or XLM-RoBERTa) experience an increase in recall but a decrease in precision, or the opposite, because the text shifts away from familiar AI phrasing. GPT+Genre further contextualizes the rewrite with domain hints (e.g. write it like a news article), causing domain-sensitive lexical changes. This can introduce new phrasing that confuses detectors reliant on domain-specific cues.
@@
XLM-RoBERTa often shows high recall (sometimes hitting 1.0), but its precision remains around 0.52, leading to only modest changes in F1. DistilBERT is more balanced yet remains vulnerable when its top keywords are replaced by synonyms or paraphrases. BERT-base typically holds a moderate advantage across scenarios but can be undermined in LIME-based rewrites (e.g., LIME+HSR, where BERT-base F1 falls to about 0.56). By contrast, the Ensemble approach consistently retains the highest or near-highest F1 and accuracy. For example, under LIME+GPT, the Ensemble still attains an F1 of 0.83, substantially above XGBoost’s 0.57 or BERT-base’s 0.73. Similarly, in SHAP plus GPT+Genre, the Ensemble preserves a strong F1 of about 0.79, while other models vary more dramatically. Also, We show text similarity scores (BLEU, ROUGE1, ROUGEL) to capture how extensively each rewrite modifies the original text.
@@
Table 6 presents text similarity scores (BLEU, ROUGE1, ROUGEL) under all these rewriting conditions. Looking at BLEU and ROUGE scores confirms that GPT-based methods, especially GPT+Genre, introduce the largest textual shifts. Replacements guided by SHAP+HSR or SHAP+PSR keep BLEU at or above 0.90 (0.93 and 0.97 respectively), while LIME/Random-based HSR achieve lower BLEU scores ( 0.84). GPT-based rewrites show BLEU scores ranging from 0.76 to 0.86 depending on the explainability method used. This trade-off highlights how more extensive rewriting can more thoroughly remove AI signatures at the expense of text fidelity.
@@
NoRewriting Original 1.00 1.00 1.00
SHAP HSR 0.93 (-7 pp) 0.97 (-3 pp) 0.97 (-3 pp)
SHAP PSR 0.97 (-3 pp) 0.99 (-1 pp) 0.99 (-1 pp)
SHAP GPT 0.86 (-14 pp) 0.93 (-7 pp) 0.93 (-7 pp)
SHAP GPT+Genre 0.86 (-14 pp) 0.94 (-6 pp) 0.94 (-6 pp)
LIME HSR 0.84 (-16 pp) 0.92 (-8 pp) 0.92 (-8 pp)
LIME PSR 0.90 (-10 pp) 0.96 (-5 pp) 0.95 (-5 pp)
LIME GPT 0.76 (-24 pp) 0.84 (-16 pp) 0.84 (-16 pp)
LIME GPT+Genre 0.76 (-24 pp) 0.85 (-15 pp) 0.84 (-16 pp)
Random HSR 0.84 (-16 pp) 0.92 (-8 pp) 0.92 (-8 pp)
Random PSR 0.90 (-11 pp) 0.95 (-5 pp) 0.95 (-5 pp)
Random GPT 0.76 (-24 pp) 0.84 (-16 pp) 0.84 (-16 pp)
Random GPT+Genre 0.76 (-24 pp) 0.84 (-16 pp) 0.84 (-16 pp)
@@
BERT $120$ 77 $5.83$
DistilBERT $120$ 99 $7.50$
XLM-R $120$ 11 $0.83$
Ensemble $120$ 88 $6.67$
PSR XGBoost $120$ 77 $5.83$
BERT $120$ 44 $3.33$
DistilBERT $120$ 44 $3.33$
XLM-R $120$ 0 $0.00$
Ensemble $120$ 33 $2.50$
@@
BERT $120$ 2020 $16.67$
DistilBERT $120$ 1919 $15.83$
XLM-R $120$ 55 $4.17$
Ensemble $120$ 1919 $15.83$
@@
SHAP GPT XGBoost $120$ 44 $3.33$
BERT $120$ 77 $5.83$
DistilBERT $120$ 77 $5.83$
XLM-R $120$ 0 $0.00$
@@
GPT+Genre XGBoost $120$ 44 $3.33$
BERT $120$ 33 $2.50$
DistilBERT $120$ 66 $5.00$
XLM-R $120$ 0 $0.00$
@@
BERT $120$ 44 $3.33$
DistilBERT $120$ 77 $5.83$
XLM-R $120$ 0 $0.00$
Ensemble $120$ 99 $7.50$
@@
BERT $120$ 66 $5.00$
DistilBERT $120$ 66 $5.00$
XLM-R $120$ 0 $0.00$
Ensemble $120$ 55 $4.17$
@@
As shown in Table 7, many rewriting strategies cause a lot of flips for some detectors, especially the simpler ones. However, the Ensemble model usually has a lower flip rate. This shows that using several models together makes detection more reliable.
Interpreting Overlap Patterns.
Figure 2 visualizes the overlap patterns among flipped samples across all models using UpSet plots. Each bar represents the number of AI-generated samples that were simultaneously flipped (misclassified as human) by a specific combination of models, with the connected dots below indicating which models are involved in that intersection.
Several patterns emerge from this analysis. First, XGBoost-only flips dominate many scenarios, particularly under LIME+HSR where XGBoost uniquely flips 40+ samples that other models correctly classify. This indicates that XGBoost’s feature-based approach is most susceptible to targeted token replacements. Second, the Ensemble exhibits notably smaller exclusive flip counts compared to other models—typically 3–11 samples depending on the scenario—demonstrating that samples fooling the Ensemble are usually also fooling other models, suggesting these are inherently difficult cases rather than ensemble-specific weaknesses.
Third, the overlap between XGBoost and transformer models is relatively limited, confirming that their failure modes differ substantially. This complementarity is precisely why the ensemble architecture succeeds: when one detection pathway fails, others often succeed. For instance, under SHAP+GPT, fewer than 5 samples fool all five models simultaneously, meaning the vast majority of evasion attempts fail against at least one component.
Fourth, XLM-RoBERTa shows consistently minimal participation in flip intersections due to its tendency to overpredict AI, which paradoxically makes it resistant to evasion attacks while reducing precision. The Ensemble balances this trade-off by combining XLM-RoBERTa’s robustness with other models’ higher precision.
These overlap patterns quantify the ensemble’s robustness advantage: even when 27.5% of samples flip under LIME+HSR for the Ensemble, the shared samples with other models suggest these represent the hardest adversarial cases where extensive token modifications fundamentally alter the text’s detectable characteristics